OpenStudy (anonymous):

A grocer sells oranges for $0.95 each and grapefruits for $1.05 each. You purchased a mix of 20 oranges and grapefruits and paid $20.20. How many of each type of fruit did you buy? (If there is no solution, enter NO SOLUTION.)

OpenStudy (klimenkov):

x - oranges y - grapefruits \(0.95x+1.05y=20.20\) \(x+y=20\) Solution is x = 8 y = 12

OpenStudy (anonymous):

HOw did you come up with the x+y=20? That is where I get hung up

OpenStudy (klimenkov):

The total number of x oranges and y grapefruits is 20. So, x+y=20.

OpenStudy (anonymous):

OMG I think I am retarded. I kept inserting 20.20! Thanks for your help, I appreciate it.

OpenStudy (klimenkov):

It's OK. Don't give up math and you will be lucky.

OpenStudy (anonymous):

I should have you handy all of the time! My test is going to kill me! lol